Missouri Journal of Mathematical Sciences Index

Problems Solved

Ahuja, Mangho, N. J. Kuenzi and Russell

Euler, Vol. 5, 46. Allard, Charles J., Vol. 1, No. 1, 34. Allard, Charles J., Vol. 1, No. 2, 32. Allard, Charles J. and Enis Alpakin, Vol.

1, No. 3, 44. Alpakin, Enis see Allard, Vol. 1, No. 3. Ashbaugh, Mark, Vol. 2, 41. Ashbaugh, Mark, Vol. 2, 146. Azarian, Mohammad K. and Jayanthi

Ganapathy, Vol. 5, 45. Azarian, Mohammad K. and J. Sriskan-

darajah, Vol. 5, 94. Azarian, Mohammad K., Vol. 3, 95. Azarian, Mohammad K., Vol. 3, 151. Azarian, Mohammad K., Vol. 5, 92. Azarian, Mohammad K., Vol. 5, 135. Bang, Seung-Jin, Vol. 4, 147. Beltramo, Alvin, Vol. 5, 98. Beltramo, Alvin, Vol. 5, 140. Chance, Joseph E., Vol. 2, 85. Chance, Joseph E., Vol. 2, 87. Chance, Joseph E., Vol. 4, 97. Chance, Joseph E., Alejandro Necochea,

Leonard L. Palmer, W. F. Wheatley, and Russell Euler, Vol. 1, No. 3, 45.

Chance, Joseph E. and Donald P. Skow,

Vol. 5, 95. Chance, Joseph E. and Jayanthi Ganapa-

thy, Vol. 4, 41. Chance, Joseph E. see Prielipp, Vol. 5, 94. Chen, Jin see Woods, Vol. 5, 147. Chen, Jin see Woods, Vol. 5, 148. Cooper, Curtis and Robert E. Kennedy,

Vol. 2, 37. Cooper, Curtis and Robert E. Kennedy,

Vol. 2, 82. Cooper, Curtis and Robert E. Kennedy,

Vol. 3, 97. Cooper, Curtis and Robert E. Kennedy,

Vol. 4, 95.

Cooper, Curtis see Doucette, Vol. 4, 149. Cooper, Curtis see Kennedy, Vol. 1, No.

3. Dixon, Albert, Vol. 1, No. 1, 33. Doucette, Robert L., Vol. 3, 47. Doucette, Robert L., Vol. 4, 30. Doucette, Robert L., Vol. 4, 31. Doucette, Robert L., Vol. 4, 34. Doucette, Robert L., Vol. 4, 39. Doucette, Robert L., Kandasamy Muthu-

vel, Curtis Cooper, and Robert E. Kennedy, Vol. 4, 149.

Doucette, Robert L. see Prielipp, Vol. 5,

94. Doucette, Robert L. see Tao, Vol. 5. Edmundson, Richard see Euler, Vol. 4, 93. Eifler, Larry, Vol. 1, No. 2, 34. Euler, Russell and James H. Taylor, Vol.

2, 88. Euler, Russell, Vol. 1, No. 1, 32. Euler, Russell, Vol. 3, 156. Euler, Russell, Vol. 4, 33. Euler, Russell, Vol. 5, 47. Euler, Russell and Joe Howard, Vol. 4, 94. Euler, Russell and Kandasamy Muthuvel,

Vol. 4, 92. Euler, Russell and Richard Edmundson,

Vol. 4, 93. Euler, Russell see Ahuja, Vol. 5. Euler, Russell see Chance, Vol. 1, No. 3. Euler, Russell see Necochea, Vol. 4. Euler, Russell see Prielipp, Vol. 2, 144. Euler, Russell see Prielipp, Vol. 5, 94. Flanigan, Frank, Vol. 5, 96. Flanigan, Frank see Prielipp, Vol. 5, 94. Flowers, Joe and Dale Woods, Vol. 1, No.

2, 31. Flowers, Joe see Witt, Vol. 3. Freitag, Hertag T. see Prielipp, Vol. 5, 94. Ganapathy, Jayanthi, Vol. 3, 49. Ganapathy, Jayanthi, Vol. 3, 153. Ganapathy, Jayanthi see Azarian, Vol. 5.

Volume 6, 1994 – Volume 10, 1998

Ganapathy, Jayanthi see Chance, Vol. 4. Gunawardena, K. L. D., Vol. 5, 98. Gunawardena, K. L. D. see Tao, Vol. 5. Hicks, Troy, Vol. 3, 51. Hicks, Troy, Vol. 4, 36. Horner, James, Vol. 1, No. 2, 35. Horner, James, Vol. 1, No. 2, 35. Howard, Joe see Euler, Vol. 4, 94. Kennedy, Robert E. and Curtis Cooper,

Vol. 1, No. 3, 42. Kennedy, Robert E. see Cooper, Vol. 2,

37. Kennedy, Robert E. see Cooper, Vol. 2,

82. Kennedy, Robert E. see Cooper, Vol. 3. Kennedy, Robert E. see Cooper, Vol. 4. Kennedy, Robert E. see Doucette, Vol. 4,

149. Kuenzi, N. J., Vol. 5, 144. Kuenzi, N. J. and Kandasamy Muthuvel,

Vol. 5, 41. Kuenzi, N. J. see Ahuja, Vol. 5. Kuenzi, N. J. see Tao, Vol. 5. Kuenzi, N. J. see Widmer, Vol. 5. Muthuvel, Kandasamy, Vol. 3, 154. Muthuvel, Kandasamy, Vol. 4, 40. Muthuvel, Kandasamy, Vol. 5, 48. Muthuvel, Kandasamy see Doucette, Vol.

4, 149. Muthuvel, Kandasamy see Euler, Vol. 4,

92. Muthuvel, Kandasamy see Kuenzi, Vol. 5. Muthuvel, Kandasamy see Prielipp, Vol. 5,

94. Necochea, Alejandro, Russell Euler, and

Miguel Paredes, Vol. 4. 96. Necochea, Alejandro see Chance, Vol. 1,

No. 3. Necochea, Alejandro see Prielipp, Vol. 2,

144. Palmer, Leonard L., Vol. 4, 38.

Palmer, Leonard L. see Chance, Vol. 1,

No. 3. Paredes, Miguel see Necochea, Vol. 4. Prielipp, Bob, Russell Euler, and Alex

Necochea, Vol. 2, 144. Prielipp, Bob, Vol. 1, No. 3, 44. Prielipp, Bob, Vol. 2, 39. Prielipp, Bob, Vol. 3, 96. Prielipp, Bob, Vol. 4, 90. Prielipp, Bob, Vol. 4, 91. Prielipp, Bob, Vol. 4, 91. Prielipp, Bob, Vol. 5, 133. Prielipp, Bob, Robert L. Doucette, Kevin

Robinson, Joseph E. Chance, Frank Flanigan, Kandasamy Muthuvel, Hertag T. Freitag, J. Sriskandarajah, and Russell Euler, Vol. 5, 94.

Rabinowitz, Stanley, Vol. 2, 84. Rabinowitz, Stanley, Vol. 3, 46. Rabinowitz, Stanley, Vol. 3, 50. Rabinowitz, Stanley, Vol. 3, 95. Rabinowitz, Stanley, Vol. 4, 148. Redmond, Don, Vol. 3, 48. Redmond, Don, Vol. 3, 155. Robinson, Kevin, Vol. 5, 93. Robinson, Kevin see Prielipp, Vol. 5, 94. Robinson, Kevin see Tao, Vol. 5. Rothbart, Andrea, Vol. 4, 147. Skow, Donald P. see Chance, Vol. 5. Sriskandarajah, J. see Azarian, Vol. 5, 94. Sriskandarajah, J. see Prielipp, Vol. 5, 94. Tao, Jenhua, Robert L. Doucette, K. L. D.

Gunawardena, N. J. Kuenzi, and Kevin Robinson, Vol. 5, 97.

Taylor, James H., Vol. 2, 86. Taylor, James H., Vol. 4, 99. Taylor, James H. see Euler, Vol. 2, 88. Thuong, Tran van, Vol. 2, 143. Wagon, Stan, Vol. 4, 150. Wheatley, W. F. see Chance, Vol. 1, No.

3.

Widmer, Lamarr and N. J. Kuenzi, Vol. 5,

136. Witt, Ana and Joe Flowers, Vol. 3, 99. Woods, Dale and Jin Chen, Vol. 5, 147. Woods, Dale and Jin Chen, Vol. 5, 148. Woods, Dale see Flowers, Vol. 1, No. 2. Ahuja, Mangho, N. J. Kuenzi and Russell

Euler, Vol. 5, 46. Allard, Charles J., Vol. 1, No. 1, 34. Allard, Charles J., Vol. 1, No. 2, 32. Allard, Charles J. and Enis Alpakin, Vol.

1, No. 3, 44. Alpakin, Enis see Allard, Vol. 1, No. 3. Ashbaugh, Mark, Vol. 2, 41. Ashbaugh, Mark, Vol. 2, 146. Azarian, Mohammad K. and Jayanthi

Ganapathy, Vol. 5, 45. Azarian, Mohammad K. and J. Sriskan-

darajah, Vol. 5, 94. Azarian, Mohammad K., Vol. 3, 95. Azarian, Mohammad K., Vol. 3, 151. Azarian, Mohammad K., Vol. 5, 92. Azarian, Mohammad K., Vol. 5, 135. Bang, Seung-Jin, Vol. 4, 147. Beltramo, Alvin, Vol. 5, 98. Beltramo, Alvin, Vol. 5, 140. Chance, Joseph E., Vol. 2, 85. Chance, Joseph E., Vol. 2, 87. Chance, Joseph E., Vol. 4, 97. Chance, Joseph E., Alejandro Necochea,

Leonard L. Palmer, W. F. Wheatley, and Russell Euler, Vol. 1, No. 3, 45.

Chance, Joseph E. and Donald P. Skow,

Vol. 5, 95. Chance, Joseph E. and Jayanthi Ganapa-

thy, Vol. 4, 41. Chance, Joseph E. see Prielipp, Vol. 5, 94. Chen, Jin see Woods, Vol. 5, 147. Chen, Jin see Woods, Vol. 5, 148. Cooper, Curtis and Robert E. Kennedy,

Vol. 2, 37.

Cooper, Curtis and Robert E. Kennedy,

Vol. 2, 82. Cooper, Curtis and Robert E. Kennedy,

Vol. 3, 97. Cooper, Curtis and Robert E. Kennedy,

Vol. 4, 95. Cooper, Curtis see Doucette, Vol. 4, 149. Cooper, Curtis see Kennedy, Vol. 1, No.

3. Dixon, Albert, Vol. 1, No. 1, 33. Doucette, Robert L., Vol. 3, 47. Doucette, Robert L., Vol. 4, 30. Doucette, Robert L., Vol. 4, 31. Doucette, Robert L., Vol. 4, 34. Doucette, Robert L., Vol. 4, 39. Doucette, Robert L., Kandasamy Muthuvel, Curtis Cooper, and Robert E. Kennedy, Vol. 4, 149.

Doucette, Robert L. see Prielipp, Vol. 5,

94. Doucette, Robert L. see Tao, Vol. 5. Edmundson, Richard see Euler, Vol. 4, 93. Eifler, Larry, Vol. 1, No. 2, 34. Euler, Russell and James H. Taylor, Vol.

2, 88. Euler, Russell, Vol. 1, No. 1, 32. Euler, Russell, Vol. 3, 156. Euler, Russell, Vol. 4, 33. Euler, Russell, Vol. 5, 47. Euler, Russell and Joe Howard, Vol. 4, 94. Euler, Russell and Kandasamy Muthuvel,

Vol. 4, 92. Euler, Russell and Richard Edmundson,

Vol. 4, 93. Euler, Russell see Ahuja, Vol. 5. Euler, Russell see Chance, Vol. 1, No. 3. Euler, Russell see Necochea, Vol. 4. Euler, Russell see Prielipp, Vol. 2, 144. Euler, Russell see Prielipp, Vol. 5, 94. Flanigan, Frank, Vol. 5, 96. Flanigan, Frank see Prielipp, Vol. 5, 94.